Complete the square as an aid in graphing:
y = -x^2 - 4x - 6
Please show all work.

Answer by lwsshak3(11628)

Complete the square as an aid in graphing:
y= -x^2 - 4x - 6
make sure coefficient of x^2=1 (in this case, it is already 1)
y=-(x^2-4x+4)-4-6
take half the coefficient of the x-term(4/2=2
square it(2^2=4
add 4 within the parenthesis
add-4 outside the parenthesis so as not to change the equation
y=-(x^2-4x+4)-4-6
y=-(x-2)^2)-10
factor the perfect square within the parenthesis, taking the sign of the x-term
combine like terms outside the parenthesis.